[figure: 9.7 cm]

MAGNETIC PROPERTIES AND MAGNETORESISTANCE OF

THE DyMnB_{4}

E.M. Levin^{1,3}, T. Palewski^{1}, T. Mydlarz^{1}, B.S. Kuzhel^{2}, Yu.B. Kuz’ma^{2} and S.I. Miklashenko^{2} 1International Laboratory of High Magnetic Fields and Low Temperature, 95 Gajowicka Str., 53-421 Wroclaw, Poland. 2Lviv State University, 1 Universitets’ka Str., Lviv, 290602 Ukraine 3Ames laboratory, Iowa State University, Ames, IA, 50011, USA

The magnetization of the DyMnB_{4} compound at 4.2 K in a magnetic field up to 13 T shows the presence of two nonlinear parts, one in magnetic field up to 4 T, and second between 4 and 13 T. Magnetization at 4.2 K in a field 13 T is equal 130 emu/g what is correspond to 6.7 μ_{B}/FU. In the temperature range 80...200 K temperature dependence of susceptibility is following C-W law with Θ = 30 K and effective magnetic moment p_{eff}= 7.5 μ_{B} what is significantly less than the theoretical value of 10.65 μ_{B} for Dy^{3+} ion. The temperature dependence of the resistance in a magnetic field shows a large magnetic contribution below 20 K and Kondo-like behaviour. The magnetoresistance (MR=[ρ(H, T)-ρ(0, T)]/ρ(0, T)) is about few percent but below 20 K increases very strongly and reaches value of 300 % at 4.2 K. A large MR for the intermetallic compound of rare-earth and transition metal is rather unusual and very probably is due to a presence of ferromagnetic order in this compound.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : Tomasz Palewski

Address for correspondence : International Laboratory of High Magnetic Fields and Low Temperature, 95 Gajowicka Str., 53-421 Wroclaw, Poland.

Email address : paltom@alpha.mlspmint.pan.wroc.pl